OpenStudy (anonymous):

Solve the equation. 6b + 4 = 2(2b + 1)

OpenStudy (anonymous):

6b+4=4b+2 2b+4=2 2b=-2 b=-1

OpenStudy (anonymous):

6b-4b=2-4 2b=-2 b=-1